A small ball of mass 1 kg is inside a circle of radius 2m at one end of the diameter. If the small ball covers quarter of the circle the displacement of the circle of radius 2m will be? Mass of large circle is 10 kg ?

Grade:11

1 Answers

Eshan
5 years ago
Dear student,

It seems as if the problem is from the center of mass concept, however some more information must have been given about the environment so that we could safely assume that the center of mass of the system does not move.

Assuming so, let the displcament of the circle bexin horizontal direction. Hence displacement of small ball will beR-x. Since center of mass does not move,Mx=m(R-x)
\implies 10x=R-x \implies x=\dfrac{R}{11}=\dfrac{2}{11}m
Similary displacement of circle in vertical direction=y=\dfrac{2}{11}m
Hence net displacment of the circle=\dfrac{2\sqrt{2}}{11}m